In a recent awards​ ceremony, the age of the winner for best actor was 32 and the age of the winner for best actress was 50. For all best​ actors, the mean age is 41.3 years and the standard deviation is 5.4 years. For all best​ actresses, the mean age is 35.3 years and the standard deviation is 10.2 years.​ (All ages are determined at the time of the awards​ ceremony.) Relative to their​ genders, who had the more extreme age when winning the​ award, the actor or the​ actress? Explain.

Since the z score for the actor is z=? and the z score for the actress is z=?​, the actress/actor had the more extreme age.
1
Expert's answer
2020-11-23T18:50:38-0500

"\bar x_{actor}=41.3"

"s_{actor}=5.4"

"x_{actor}=32"

"\bar x_{actress}=35.3"

"s_{actress}=10.2"

"x_{actress}=50"


"Z=\frac{x-\bar x}{s}"

So

"Z_{actor}=\frac{32-41.3}{5.4}=-1.72"

"Z_{actress}=\frac{50-35.3}{10.2}=1.44"


Since the value of "Z_{actor}" is higher than the value of "Z_{actress}" (absolute value), thus the value of "x_{actor}" is further from the mean compare to the value of "x_{actress}". Therefore the actor has more extreme age.